OpenStudy (anonymous):

Find the 7th term of the geometric sequence with the given terms. a4 = -8 and a5 = -40; a7 = _______. help??? D:

OpenStudy (cwrw238):

7th term = ar^(7-1) where a = first term and r = common ratio

OpenStudy (cwrw238):

r = a5 / a4

OpenStudy (cwrw238):

after you find a7 = a5 * r^2

OpenStudy (cwrw238):

* after you find r a7 = a5* r^2

OpenStudy (waleed_imtiaz):

U know. a4=a1r3 and a5=a1r4 Divide both the equations..... a5/a4 = a1r4/a1r3 -40/-8=r r=5 Put r in any of the equations to get a1... now U can find...a7=a1r6